# How Good is Post-Hoc Watermarking With Language Model Rephrasing?

### v1

@@ Abstract
Generation-time text watermarking embeds statistical signals into text for traceability of AI-generated content. We explore post-hoc watermarking where an LLM rewrites existing text while applying generation-time watermarking, to protect copyrighted documents, or detect their use in training or RAG via watermark radioactivity. Unlike generation-time approaches, which is constrained by how LLMs are served, this setting offers additional degrees of freedom for both generation and detection. We investigate how allocating compute (through larger rephrasing models, beam search, multi-candidate generation, or entropy filtering at detection) affects the quality-detectability trade-off. Our strategies achieve strong detectability and semantic fidelity on open-ended text such as books. Among our findings, the simple Gumbel-max scheme surprisingly outperforms more recent alternatives under nucleus sampling, and most methods benefit significantly from beam search. However, most approaches struggle when watermarking verifiable text such as code, where we counterintuitively find that smaller models outperform larger ones. This study reveals both the potential and limitations of post-hoc watermarking, laying groundwork for practical applications and future research.
@@ 1 Introduction
Post-hoc text watermarking inserts an algorithmically-detectable signal into an existing text while remaining imperceptible to readers. It serves several goals such as copyright protection or traitor tracing. Early approaches relied on hand-crafted modifications: synonym substitutions (Topkara et al., 2006c; Shirali-Shahreza and Shirali-Shahreza, 2008), grammatical transformations (Topkara et al., 2006b, a) or morphosyntactic alterations (Meral et al., 2009). Recent methods based on deep neural networks follow a post-hoc paradigm where the input is the original text and the output is the watermarked text, using one watermark embedder and one watermark extractor (Abdelnabi and Fritz, 2021), similarly to state-of-the-art post-hoc watermarking approaches for images, audio, and video. However, these methods are not effective, suffering from low capacity or unreliable detection, because they are highly constrained or easily broken by reversing the edits or synonyms.
@@
With the emergence of large language models (LLMs), their popularization through ChatGPT (OpenAI, 2022), and growing concerns about potential risks (Crothers et al., 2022; Weidinger et al., 2022), generation-time text watermarking algorithms have been introduced to help detect AI-generated text (Aaronson and Kirchner, 2023; Kirchenbauer et al., 2023a). These algorithms have been deployed at scale, for instance in Google’s Gemini with SynthID (Dathathri et al., 2024). Most LLM watermarks alter the next token selection, for example, by promoting a specific set of tokens depending on previous tokens and a secret key. Detection of the watermark in a text is then performed through a theoretically grounded statistical test that provides rigorous guarantees over the false positive rates. In addition, these methods achieve high watermark power while adding minimal latency, by leveraging the entropy of the LLM to embed the signal. However, in contrast with previous approaches, these methods require control at generation time and therefore cannot be applied to existing text.
@@
A natural idea to extend these methods to post-hoc watermarking is to employ a paraphrasing LLM to re-generate the text, allowing the watermark to be injected at inference time without further model training. This approach has been explored in data protection literature (Jovanović et al., 2025; Sander et al., 2025; Rastogi et al., 2025; Zhang et al., 2025; Lau et al., 2024). For example, “watermark radioactivity” exploits the fact that watermark signals leave detectable traces when watermarked text is used by another model, enabling active training and context membership inference (Sander et al., 2025; Jovanović et al., 2025). However, to our knowledge, there is no thorough evaluation of how post-hoc watermarking through LLM paraphrasing performs across different data domains, such as prose versus verifiable text like code. Moreover, unlike generation-time watermarking where watermark embedding should not delay the generation, post-hoc watermarking allows trading additional compute for a better quality-detectability trade-off. This opens up several axes to explore: model size, watermarking method, decoding strategy (and even generating multiple candidates), as well as detection strategy.
@@ 2.2 Generation-Time Watermarking With Large Language Models
Extensions include improved tests and multi-bit watermarking (Fernandez et al., 2023; Yoo et al., 2023b, 2024; Qu et al., 2024), position-dependent seeds (Christ et al., 2023; Kuditipudi et al., 2023), low-entropy optimizations (Lee et al., 2023; Christ et al., 2023; Huang et al., 2023), and semantic watermarks for better robustness (Liu et al., 2023; Liu and Bu, 2024; Fu et al., 2024; Hou et al., 2023, 2024). DiPMark (Wu et al., 2023) provides distortion-free Green-Red watermarks, and MorphMark (Wang et al., 2025) adaptively adjusts watermark strength based on the green token probability mass. WaterMax (Giboulot and Furon, 2024) generates several chunks of tokens from the original LLM distribution and selects the outputs with high watermark scores, which ensures that the LLM distribution is preserved. SynthID-Text (Dathathri et al., 2024) deploys tournament-based sampling in Google Gemini. Toolkits have also been introduced to benchmark these methods (Piet et al., 2023; Pan et al., 2024).
@@ 2.3 Post-Hoc LLM Watermarks for Data Protection
Recent works apply LLM watermarks to training or evaluation data via paraphrasing, similarly as what we study in this work. They exploit watermark radioactivity (Sander et al., 2024), i.e., the detectable traces left when watermarked text is used for training. Applications include detection of texts used in retrieval augmented generation (RAG) (Jovanović et al., 2025), benchmark contamination detection (Sander et al., 2025), and training data copyright (Zhang et al., 2025). Only Waterfall (Lau et al., 2024) evaluates post-hoc watermarking through LLM paraphrasing, focusing on code (MBPP) and natural text (C4 and arXiv) for provenance detection.
These works demonstrate the utility of such method for data protection but do not evaluate it as a general watermarking one and do not characterize failure modes across text types and settings.
@@ 3.1 LLM Watermarking
Detection recomputes the pseudorandom function for each group of tokens and aggregates the results into a test statistic. Under the null hypothesis $\mathcal{H}_{0}$ “not watermarked with this specific scheme”, the statistic should follow a known distribution, yielding a $p$-value for the probability of observing such a score or a bigger score by chance. Under the alternative hypothesis $\mathcal{H}_{1}$ “watermarked text with this specific scheme”, the statistic should deviate significantly enough and lead to low $p$-values. We flag text as watermarked if the $p$-value is below a detection threshold $\alpha$, which controls the false positive rate (FPR).
For instance, if $K$ (the test statistic here) is the observed number of green tokens out of $N$ scored tokens and that this statistic follows under $\mathcal{H}_{0}$ a Binomial distribution with parameters $N$ and $\gamma$ (expected green ratio, typically $0.5$), then $p\text{-value}=\mathbb{P}\left(X\geq K\mid X\sim\text{Binomial}(N,\gamma)\right)$. As an example, if we observe $K=65$ green tokens out of $N=100$ with $\gamma=0.5$, the $p$-value is approximately $10^{-3}$; to flag this text as watermarked, we would need a FPR of $\alpha\geq 10^{-3}$.
@@ 3.2 Compute-Driven Flexibility
We use a beam search approach to improve the quality-detectability trade-off. Specifically, we maintain $B$ candidate sequences and expands each beam with $V$ candidates sampled from the watermarked probability distribution $\mathbf{p}_{\text{wm}}(\cdot\mid x_{<t})$, which is derived by applying the watermarking scheme (e.g., Maryland, SynthID, DIPMark, MorphMark) to the base model’s logits. To select the top-$B$ beams at each step, we score candidates using log-probability under a reference distribution. We explore two scoring variants: unbiased scoring uses the original model probabilities $\mathbf{p}_{\text{orig}}(\cdot\mid x_{<t})$ to favor sequences with lower perplexity relative to the base model, prioritizing text quality; biased scoring uses the watermarked probabilities $\mathbf{p}_{\text{wm}}(\cdot\mid x_{<t})$ to favor sequences that are most probable under the watermarked distribution, potentially yielding stronger watermark signals. We note that the latter was also explored for generation-time watermarking in (Kirchenbauer et al., 2023b). For candidate generation, we consider both deterministic beam search (selecting top-$V$ tokens by $\mathbf{p}_{\text{wm}}$) and stochastic beam search (sampling $V$ tokens from $\mathbf{p}_{\text{wm}}$).
Multi-candidate selection with Wateremax (Giboulot and Furon, 2024)
Rephrasing is done chunk by chunk of $L$ tokens. For each chunk, we generate $m$ candidates $\{\tilde{y}^{(1)},\dots,\tilde{y}^{(m)}\}$ without applying any logit bias. We then select the candidate that naturally maximizes the watermark score: $y^{*}=\arg\max_{y}\text{Score}_{\text{wm}}(y)$. This method is “distortion-free” regarding the sampling distribution but is typically infeasible in standard API usage due to cost (generating $L$ times more tokens per chunk).
@@ 4.1 Experimental Setup
  • Books: Passages in English from the Gutenberg dataset (Project Gutenberg, 2025), containing books. These represent longer-form, literary text with complex sentence structures.
  • Wikipedia: Lead paragraphs from 500 randomly sampled Wikipedia articles. These represent factual, encyclopedic text with a formal writing style and different languages.
@@
To ensure the models outputs the watermarked text directly and without any comments, we use a structured system prompt and prefill the assistant’s response with a prefix (e.g., “Rephrased text:”). An example prompt used is:
@@
  • SynthID-Text (Dathathri et al., 2024): We vary the number of tournaments $k\in\{10,20,30\}$. We also compute the $p$-values from weighted and non weighted scores as described in subsection 6.2.
@@
We address the private key issue described in section 3 by testing 50 candidate keys. For each combination of tokenizer, watermarking method (and depth for SynthID), we evaluate the keys on non-watermarked texts to estimate the $p$-value distribution under $H_{0}$. We then select the key that minimizes the deviation from $U(0,1)$, as measured by the Kolmogorov–Smirnov statistic. Next, we verify that the theoretical FPRs for these keys match the empirical FPRs on 1.5M Wikipedia documents containing between 1k and 100k tokens (using watermark window deduplication). This ensures that the subsequent results under $H_{1}$ reflect the intrinsic performance of each scheme rather than artifacts of a particular key choice. Further details are provided in Table 6.
@@
For all methods, we evaluate entropy-aware detection by computing per-token entropy with the same model used for rephrasing. At detection, we only score tokens with entropy $\geq\tau$, with $\tau\in\{0,0.2,0.4,0.6,\dots,1.8,2.0\}$. We note that we only pass the watermarked text to compute the per-token entropy without the original text as context.
@@
  • Beam search (Sutskever et al., 2014), as detailed in section 3. We test beam widths $B\in\{3,5,10\}$ with same number of candidates per beam. We experiment with selection criteria based on either the non-watermarked likelihood or the watermarked model’s likelihood, and use stochastic (temperature $1$ and top-$p$ 0.95) or deterministic selection within beams.
@@
  • Detection: We compute the $p$-value of the corresponding detection test, which can be read as the probability that a watermark score at least as high would happen if the text was not watermarked.
  • Quality: We evaluate rephrase quality along three axes. Semantic similarity: We use the BERTScore (Zhang et al., 2020) to measure meaning preservation between original and rephrased texts. Naturalness: We compute the average cross entropy of the rephrased text using Mistral-7B-v0.3 (Jiang et al., 2023), conditioned on the prefix “This is a rephrased version of [ORIGINAL TEXT]:”, capturing if the output is fluent and plausible. Length ratio: We report the character-level length ratio between rephrased and original text, which should remain close to $1$.
@@
In Sections 4.2, 4.3, 4.5, 4.4 and 4.8, each reported value corresponds to the median over 100 passages of approximately 800 characters ($\sim$200 tokens) sampled from Charles Dickens’ novels in the Project Gutenberg dataset (Project Gutenberg, 2025). We strictly retain only rephrasings where the output length remains comparable to the original (length ratio $\in[0.75,1.25]$), reporting results only for settings where at least 70% of passages satisfy this constraint.
@@ 4.2 Quality-Detection Trade-Off
We first evaluate the fundamental trade-off between rephrasing quality and detection strength in Figure 2, using Llama-3.2-3B-Instruct with random sampling. Each point corresponds to a distinct parameter configuration as described in subsection 4.1. Optimal methods occupy the top-right region of the plot, combining high text quality with strong detectability. The Gumbel-max method appears to dominate this Pareto frontier, while DiPMark and MorphMark lag behind.
@@
4.3 Impact of Model Family and Scale
We investigate whether larger models are better suited for post-hoc watermarking. Figure 3 shows the trade-off curves by model family (SmolLM, Gemma, Llama, Qwen) and size. Regarding performance, we observe that larger models preserve semantics better: across all families, increasing model size (darker points) shifts the clusters upward. For example, within the Llama family, Llama-3.1-8B (dark green) attains lower cross entropy than the 3B and 1B variants (lighter teal) at comparable watermark strengths. However, larger models struggle to reach low $p$-values compared to smaller models: they populate the left rather than the right of the plots. Gemma 3 models are essentially absent from the high-detectability region, as their outputs naturally exhibit very low entropy, even for the smallest sizes. Overall, when strong watermarking is required, only small models appear on the frontier, whereas larger models dominate in the lower-strength regime. We see in subsection 4.6 that the issue with large models persists even if temperature is increased.
Note that these results are conditioned on the length filtering described in subsection 4.1; thus, the presence of a data point implies enough successful valid generation for the corresponding combination of model/scheme/parameters. We refer the reader to section 8 to assess the generation stability of each configuration, as smaller models fail to meet length constraints more frequently, cannot be read from the figure.
Figure 3: Impact of model family and size. Cross Entropy vs. watermark strength. Larger models improve quality for a given watermark strength, but small models are necessary to reach high strengths. All families are comparable, except Gemma-3 that is not suitable. Experimental details are given in Sections 4.1 and 4.3.

### v2

@@ Abstract
Generation-time text watermarking embeds statistical signals into text for traceability of AI-generated content. We explore post-hoc watermarking where an LLM rewrites existing text while applying generation-time watermarking, to protect copyrighted documents, or detect their use in training or RAG via watermark radioactivity. Unlike generation-time approaches, which are constrained by how LLMs are served, this setting offers additional degrees of freedom for both generation and detection. We investigate how allocating compute (through larger rephrasing models, beam search, multi-candidate generation, or entropy filtering at detection) affects the quality-detectability trade-off. Our strategies achieve strong detectability and semantic fidelity on open-ended text such as books. Among our findings, the simple Gumbel-max scheme surprisingly outperforms more recent alternatives under nucleus sampling, and most methods benefit significantly from beam search. However, most approaches struggle when watermarking verifiable text such as code, where we counterintuitively find that smaller models outperform larger ones. This study reveals both the potential and limitations of post-hoc watermarking, laying groundwork for practical applications and future research.
@@ 1 Introduction
Post-hoc text watermarking inserts an algorithmically-detectable signal into an existing text while remaining imperceptible to readers. It serves several goals, such as copyright protection or traitor tracing. Early approaches relied on hand-crafted modifications: synonym substitutions (Topkara et al., 2006c; Shirali-Shahreza and Shirali-Shahreza, 2008), grammatical transformations (Topkara et al., 2006b, a) or morphosyntactic alterations (Meral et al., 2009). Recent methods based on deep neural networks follow a post-hoc paradigm where the input is the original text and the output is the watermarked text, using one watermark embedder and one watermark extractor (Abdelnabi and Fritz, 2021), similarly to state-of-the-art post-hoc watermarking approaches for images, audio, and video. However, these methods are not effective, suffering from low capacity or unreliable detection, because they are highly constrained or easily broken by reversing the edits or synonyms.
@@
With the emergence of large language models (LLMs), their popularization through ChatGPT (OpenAI, 2022), and growing concerns about potential risks (Crothers et al., 2022; Weidinger et al., 2022), generation-time text watermarking algorithms have been introduced to help detect AI-generated text (Aaronson and Kirchner, 2023; Kirchenbauer et al., 2023a). These algorithms have been deployed at scale, for instance, in Google’s Gemini with SynthID (Dathathri et al., 2024). Most LLM watermarks alter the next token selection, for example, by promoting a specific set of tokens depending on previous tokens and a secret key. Detection of the watermark in a text is then performed through a theoretically grounded statistical test that provides rigorous guarantees over the false positive rates. In addition, these methods achieve high watermark power while adding minimal latency by leveraging the entropy of the LLM to embed the signal. However, in contrast with previous approaches, these methods require control at generation time and therefore cannot be applied to existing text.
@@
A natural idea to extend these methods to post-hoc watermarking is to employ a paraphrasing LLM to re-generate the text, allowing the watermark to be injected at inference time without further model training. This approach has been explored in data protection literature (Jovanović et al., 2025; Sander et al., 2025; Rastogi et al., 2025; Zhang et al., 2025; Lau et al., 2024). For example, “watermark radioactivity” exploits the fact that watermark signals leave detectable traces when watermarked text is used by another model, enabling active training and context membership inference (Sander et al., 2025; Jovanović et al., 2025). However, to our knowledge, there is no thorough evaluation of how post-hoc watermarking through LLM paraphrasing performs across different data domains, such as prose versus verifiable text like code. Moreover, unlike generation-time watermarking, where watermark embedding should not delay the generation, post-hoc watermarking allows trading additional computation for a better quality-detectability trade-off. This opens up several axes to explore: model size, watermarking method, decoding strategy (and even generating multiple candidates), as well as detection strategy.
@@ 2.2 Generation-Time Watermarking With Large Language Models
Subsequent work explores improved tests and multi-bit watermarking (Fernandez et al., 2023; Yoo et al., 2023b, 2024; Qu et al., 2024), position-dependent seeds (Christ et al., 2023; Kuditipudi et al., 2023), low-entropy optimizations (Lee et al., 2023; Christ et al., 2023; Huang et al., 2023), and semantic watermarks for better robustness (Liu et al., 2023; Liu and Bu, 2024; Fu et al., 2024; Hou et al., 2023, 2024). DiPMark (Wu et al., 2023) provides distortion-free Green-Red watermarks, and MorphMark (Wang et al., 2025) adaptively adjusts watermark strength based on the green token probability mass. WaterMax (Giboulot and Furon, 2024) generates several chunks of tokens from the original LLM distribution and selects outputs with high watermark scores, which ensures that the original LLM distribution is preserved. SynthID-Text (Dathathri et al., 2024) deploys tournament-based sampling in Google Gemini. Toolkits have also been introduced to benchmark these methods (Piet et al., 2023; Pan et al., 2024).
@@ 2.3 Post-Hoc LLM Watermarks for Data Protection
Recent works apply LLM watermarks to training or evaluation data via paraphrasing, similar to what we study in this work. They exploit watermark radioactivity (Sander et al., 2024), i.e., the detectable traces left when watermarked text is used for training. Applications include detection of texts used in retrieval-augmented generation (RAG) (Jovanović et al., 2025), benchmark contamination detection (Sander et al., 2025), and training data copyright (Zhang et al., 2025). Only Waterfall (Lau et al., 2024) evaluates post-hoc watermarking through LLM paraphrasing, focusing on code (MBPP) and natural text (C4 and arXiv) for provenance detection.
These works demonstrate the utility of such methods for data protection, but do not evaluate it as a general watermarking method, nor do they characterize failure modes across text types and settings.
@@ 3.1 LLM Watermarking
Detection recomputes the pseudorandom function for each group of tokens and aggregates the results into a test statistic. Under the null hypothesis $\mathcal{H}_{0}$ “not watermarked with this specific scheme”, the statistic should follow a known distribution, yielding a $p$-value for the probability of observing such a score or a more extreme score by chance. Under the alternative hypothesis $\mathcal{H}_{1}$ “watermarked text with this specific scheme”, the statistic should deviate significantly enough and lead to low $p$-values. We flag text as watermarked if the $p$-value is below a detection threshold $\alpha$, which controls the false positive rate (FPR).
For instance, if $K$ (the test statistic here) is the observed number of green tokens out of $N$ scored tokens and this statistic follows under $\mathcal{H}_{0}$ a Binomial distribution with parameters $N$ and $\gamma$ (expected green ratio, typically $0.5$), then $p\text{-value}=\mathbb{P}\left(X\geq K\mid X\sim\text{Binomial}(N,\gamma)\right)$. As an example, if we observe $K=65$ green tokens out of $N=100$ with $\gamma=0.5$, the $p$-value is approximately $10^{-3}$; to flag this text as watermarked, we would need a FPR of $\alpha\geq 10^{-3}$.
@@ 3.2 Compute-Driven Flexibility
We use a beam search approach to improve the quality-detectability trade-off. Specifically, we maintain $B$ candidate sequences and expand each beam with $V$ candidates sampled from the watermarked probability distribution $\mathbf{p}_{\text{wm}}(\cdot\mid x_{<t})$, which is derived by applying the watermarking scheme (e.g., Maryland, SynthID, DIPMark, MorphMark) to the base model’s logits. To select the top-$B$ beams at each step, we score candidates using log-probability under a reference distribution. We explore two scoring variants: unbiased scoring uses the original model probabilities $\mathbf{p}_{\text{orig}}(\cdot\mid x_{<t})$ to favor sequences with lower perplexity relative to the base model, prioritizing text quality; biased scoring uses the watermarked probabilities $\mathbf{p}_{\text{wm}}(\cdot\mid x_{<t})$ to favor sequences that are most probable under the watermarked distribution, potentially yielding stronger watermark signals. We note that the latter was also explored for generation-time watermarking in (Kirchenbauer et al., 2023b). For candidate generation, we consider both deterministic beam search (selecting top-$V$ tokens by $\mathbf{p}_{\text{wm}}$) and stochastic beam search (sampling $V$ tokens from $\mathbf{p}_{\text{wm}}$).
Multi-candidate selection with WaterMax (Giboulot and Furon, 2024).
Rephrasing is done chunk by chunk, with chunks of $L$ tokens. For each chunk, we generate $m$ candidates $\{\tilde{y}^{(1)},\dots,\tilde{y}^{(m)}\}$ without applying any logit bias. We then select the candidate that naturally maximizes the watermark score: $y^{*}=\arg\max_{y}\text{Score}_{\text{wm}}(y)$. This method is “distortion-free” regarding the sampling distribution but is typically infeasible in standard API usage due to cost (generating $L$ times more tokens per chunk).
@@ 4.1 Experimental Setup
  • Books: Passages in English from the Gutenberg dataset (Project Gutenberg, 2025), containing books. These represent longer-form, literary texts with complex sentence structures.
  • Wikipedia: The beginning of randomly sampled Wikipedia articles. These represent factual, encyclopedic texts with a formal writing style and different languages.
@@
To ensure the model outputs the watermarked text directly and without any comments, we use a structured system prompt and prefill the assistant’s response with a prefix (e.g., “Rephrased text:”). An example prompt used is:
@@
  • SynthID-Text (Dathathri et al., 2024): We vary the number of tournaments $k\in\{10,20,30\}$. We also compute the $p$-values from weighted and non-weighted scores as described in subsection 6.2.
@@
We address the private key issue described in section 3 by testing 50 candidate keys. For each combination of tokenizer, watermarking method (and depth for SynthID), we evaluate the keys on non-watermarked texts to estimate the $p$-value distribution under $H_{0}$. We then select the key that minimizes the deviation from $U(0,1)$, as measured by the Kolmogorov–Smirnov statistic. Next, we verify that the theoretical FPRs for these keys match the empirical FPRs on 1.5M Wikipedia documents (they contain up to 100k tokens, with an average of 715 tokens). This ensures that the subsequent results under $H_{1}$ reflect the intrinsic performance of each scheme rather than artifacts of a particular key choice. Further details are provided in Table 6.
@@
For all methods, we evaluate entropy-aware detection by computing per-token entropy with the same model as used for rephrasing. At detection, we only score tokens with entropy $\geq\tau$, with $\tau\in\{0,0.2,0.4,0.6,\dots,1.8,2.0\}$. We note that we only pass the watermarked text to compute the per-token entropy without the original text as context.
@@
  • Beam search (Sutskever et al., 2014), as detailed in section 3. We test beam widths $B\in\{3,5,10\}$ with the same number of candidates per beam. We experiment with selection criteria based on either the non-watermarked likelihood or the watermarked model’s likelihood, and use stochastic (temperature $1$ and top-$p$ 0.95) or deterministic selection within beams.
@@
  • Detection: We compute the $p$-value of the corresponding detection test, which can be read as the probability that a watermark score at least as high would happen if the text were not watermarked.
  • Quality: We evaluate rephrasing quality along three axes. Semantic similarity: We use the BERTScore (Zhang et al., 2020) to measure meaning preservation between original and rephrased texts. Naturalness: We compute the average cross entropy of the rephrased text using Mistral-7B-v0.3 (Jiang et al., 2023), conditioned on the prefix “This is a rephrased version of [ORIGINAL TEXT]:”, capturing if the output is fluent and plausible. Length ratio: We report the character-level length ratio between rephrased and original text, which should remain close to $1$.
@@
In Sections 4.2, 4.3, 4.5, 4.4 and 4.8, each reported value corresponds to the median over 100 passages of approximately 800 characters ($\sim$200 tokens) sampled from Charles Dickens’ novels in the Project Gutenberg dataset (Project Gutenberg, 2025). We keep only rephrasings where the output length remains comparable to the original (length ratio $\in[0.75,1.25]$), and report results only for settings where at least 70% of passages satisfy this constraint.
@@ 4.2 Quality-Detection Trade-Off
We first evaluate the fundamental trade-off between rephrasing quality and detection strength in Figure 2, using Llama-3.2-3B-Instruct with random sampling. Each point corresponds to a distinct parameter configuration as described in subsection 4.1. Optimal methods occupy the top-right region of the plot, combining high text quality with strong detectability. The Gumbel-max method appears to dominate this Pareto frontier, followed by MorphMark and SynthID.
@@
4.3 Impact of Model Family and Size
We investigate whether larger models are better suited for post-hoc watermarking. Figure 3 shows the trade-off curves by model family (SmolLM, Gemma, Llama, Qwen) and size. Regarding performance, we observe that larger models preserve semantics better: across all families, increasing model size (darker points) shifts the clusters upward. For example, within the Llama family, Llama-3.1-8B (dark green) attains lower cross entropy than the 3B and 1B variants (lighter teal) at comparable watermark strengths. However, larger models struggle to reach low $p$-values compared to smaller models: they populate the left rather than the right of the plots. Gemma 3 models are essentially absent from the high-detectability region, as their outputs naturally exhibit very low entropy, even for the smallest sizes. Overall, when strong watermarking is required, only small models appear on the frontier, whereas larger models dominate in the lower-strength regime. We see in subsection 4.6 that the issue with large models persists even if the temperature is increased.
Note that these results are conditioned on the length filtering described in subsection 4.1; thus, the presence of a data point implies enough successful valid generation for the corresponding combination of model/scheme/parameters. For instance, smaller models with stronger watermarks fail to meet length constraints more frequently. We refer the reader to section 8 to assess the generation stability of each configuration, which cannot be inferred directly from the figure.
Figure 3: Impact of model family and size. Cross Entropy vs. watermark strength. Larger models improve quality for a given watermark strength, but small models are necessary to reach high strengths. All families are comparable, except Gemma-3 which is not suitable. Experimental details are given in Sections 4.1 and 4.3.